mhanifa:

How do you understand this? A woman looks down from a hot air balloon. She sees sheep below and measures the angle of depression as 35∘. If the sheep are 125 feet from where the woman is looking down from the balloon (hypotenuse), how high off the ground is the balloon? a. 72 feet b. 102 feet c. 88 feet d. 43 feet

jhonyy9:

try draw this case

mhanifa:

To me "If the sheep are 125 feet from where the woman is looking down from the balloon (hypotenuse)" bit is confusing, it could be interpreted as vertical distance to the balloon, if the "hypotenuse" is not there

jhonyy9:

@extrinix any idea here ?

Extrinix:

Well you have an angle, \(35^\circ\) and you also have the hypotenous, \(125ft\) and you need to find the height, which is the opposite to \(35^\circ\) This is my thinking because it's not the angle shown from the balloon to the ground, it's the angle opposite of the height.

Extrinix:

So it would be \(sin(35^\circ)=\dfrac{x}{125}\)

mhanifa:



Extrinix:

x is on the wrong side, the hypotenous is 125 and the length of the `opposite of 35°` is x

mhanifa:



Extrinix:

no, 125 is the `hypotenous` only

Extrinix:

you would just leave the bottom side blank

Extrinix:

leg*

mhanifa:

"If the sheep are 125 feet from where the woman is looking down from the balloon" , suggests the horizontal leg is 125 ft, but adding the word "hypotenuse" adds confusion

Extrinix:

"if the sheep are 125 feet from where the woman is look down from the balloon (hypotenous)" Read further

Extrinix:

It's saying that the length from the woman to the sheep is your hypotenous.

Extrinix:

hypotenuse** Autocorrect to hypotenous?? Ē

mhanifa:

Thanks, this is my understanding of the problem but the thing is it leads to incorrect answer. And the correct answer choice given is c. 88 feet. This makes me to think the question is incorrect. And this is the word "hypotenuse"

mhanifa:

And my answer choice is a. 72 feet, as you suggest too

Extrinix:

Maybe it's, \(cos(90-35^\circ)=\dfrac{x}{125}\) ??

Extrinix:

No, because it would still be 72...

mhanifa:

That comes from tan 35

Extrinix:

Yeah I think either the correct answer choice is wrong or the question itself is wrong...

Extrinix:

it also comes from cos 55

mhanifa:

Thanks for your time.

Extrinix:

We tried, but I'm guessing a messup in the question itself, not our calculations.

mhanifa:

Yeah, the purpose of this question was to confirm this.

Extrinix:

Ah, okay.